### v1

# SelectIT: Selective Instruction Tuning for Large Language Models via Uncertainty-Aware Self-Reflection

Abstract
Instruction tuning (IT) is crucial to tailoring large language models (LLMs) towards human-centric interactions. Recent advancements have shown that the careful selection of a small, high-quality subset of IT data can significantly enhance the performance of LLMs. Despite this, common approaches often rely on additional models or data sets, which increases costs and limits widespread adoption. In this work, we propose a novel approach, termed SelectIT, that capitalizes on the foundational capabilities of the LLM itself. Specifically, we exploit the intrinsic uncertainty present in LLMs to more effectively select high-quality IT data, without the need for extra resources. Furthermore, we introduce a novel IT dataset, the Selective Alpaca, created by applying SelectIT to the Alpaca-GPT4 dataset. Empirical results demonstrate that IT using Selective Alpaca leads to substantial model ability enhancement. The robustness of SelectIT has also been corroborated in various foundation models and domain-specific tasks. Our findings suggest that longer and more computationally intensive IT data may serve as superior sources of IT, offering valuable insights for future research in this area. Data, code, and scripts are freely available at https://github.com/Blue-Raincoat/SelectIT.
1 Introduction
Large language models (LLMs) have attracted much attention due to their impressive capabilities in following instructions and solving intricate problems (Touvron et al., 2023b, a; Achiam et al., 2023; Penedo et al., 2023). A crucial aspect of enhancing LLMs’ performance is instruction tuning (IT), which involves the supervised adjustment of LLMs using pairs of instructional data, essential for refining the models’ ability to accurately respond to human instructions. Recent groundbreaking research, such as the LIMA Zhou et al. (2023), highlights the critical importance of instructional data quality over quantity. Contrary to the approach of merely increasing the dataset size, a carefully selected, smaller dataset of higher quality can significantly improve LLMs’ performance.
Figure 1: Existing advanced data selection strategies rely heavily on external models or data sets; however, SelectIT effectively overcomes this limitation.
Despite the emergence of various methods for selecting high-quality data, these methods often rely on external resources and limit broader implementation. External ModelChen et al. (2023a); Liu et al. (2023) propose the employment of closed-source LLMs to evaluate or rank IT data. To circumvent the closed-source limitations, Li et al. (2023a, b); Kung et al. (2023) recommend the adoption of fine-tuned open-source LLMs, which will necessitate increased computational resource expenditure. External DataCao et al. (2023) split all mixed data into several bins and fully trained the models to evaluate different indicators of high-quality IT data. Despite these advancements, the challenge of conducting precise and efficient high-quality data selection without external resources remains unresolved.
We introduce SelectIT, a novel approach designed to enhance IT data selection by fully leveraging the foundation model itself, eliminating the need for external resources. SelectIT employs different grain uncertainty of LLMs: token, sentence, and model, which can effectually improve the accuracy of IT data selection. We first use the foundation model itself to rate the IT data from 1 to $K$ based on the uncertainty of various tokens. Next, we use sentence-level uncertainty to improve the rating process by exploiting the effect of different prompts on LLMs. At a higher model level, we utilize the uncertainty between different LLMs, enabling a collaborative decision-making process for IT data selection. By applying SelectIT to the original Alpaca, we curate a compact and superior IT dataset, termed Selective Alpaca.
Experimental results show that SelectIT outperforms existing high-quality data selection methods, improving LLaMA-2’s performance on the open-instruct benchmark Wang et al. (2023). Further analysis reveals that SelectIT can effectively discard abnormal data and tends to select longer and more computationally intensive IT data. The primary contributions of SelectIT are as follows:
  • We propose SelectIT, a novel IT data selection method which exploits the uncertainty of LLMs without using additional resources.
  • We introduce a novel IT dataset, Selective Alpaca, by selecting the high-quality IT data from the Alpaca-GPT4 dataset.
  • SelectIT can substantially improve the performance of LLMs across a variety of foundation models and domain-specific tasks.
  • Our analysis suggests that longer and more computationally intensive IT data may be more effective, offering a new perspective on the characteristics of optimal IT data.
2 Related Work
Instruction Tuning Dataset
[…]
Instruction Data Selection
The recognition of IT data quality’s superiority over quantity in the context of instruction tuning is well-established, yet the efficient and precise identification of high-quality data continues to be a challenging frontier for research. One straightforward approach is utilizing the closed-source advanced LLMs for IT data evaluation and selection Chen et al. (2023a); Liu et al. (2023). To circumvent the constraints associated with closed-source, some researchers opt to fine-tune LLMs directly to select high-quality IT data Li et al. (2023b); Kung et al. (2023). Instruction Mining Cao et al. (2023), involves training multiple models to derive various metrics, using statistical regression to develop a data quality assessment framework. Li et al. (2023a) propose training a specialized model and utilizing two unique, condition-based losses on this for a comprehensive IT data selection. Additionally, Wu et al. (2023) explore where data selection is informed by the similarity of samples within the embedding space of a fine-tuned model. However, the methods described above depend, to varying degrees, on supplementary datasets, the use of closed-source models, or open-source models that have been specially fine-tuned, which results in increased consumption of resources and limits the broader impact.
Figure 2: The overall framework of SelectIT.
3 Our SelectIT Method
Utilizing advanced LLMs for the sample evaluation is a widely adopted approach in the IT data selection Chen et al. (2023a); Li et al. (2023b); Liu et al. (2023). Given an IT dataset $D$ containing a sample $S$ = (input $X$, response $Y$), a designated rating prompt $RP$, and the LLMs $M$, the goal is to leverage both $RP$ and $S$ to prompt $M$ to assign an evaluation $Score$ to the sample $S$ on a scale from 1 to $K$. A higher score typically signifies superior IT data $Quality$.
$Quality\propto Score\in[1,K]=M(RP,S)$ (1)
While existing methods Chen et al. (2023a); Cao et al. (2023) are adept at identifying high-quality samples, they often over-rely on external resources. To address these challenges, we introduce SelectIT, a strategy that capitalizes on the internal uncertainty of LLMs to efficiently select high-quality IT data. SelectIT incorporates three grains of sample evaluation modules: token, sentence, and model-level self-reflections, which effectively improve the reliability of IT data selection. The comprehensive framework of SelectIT is depicted in Figure 2.
3.1 Token-level Self-Reflection
Numerous studies have demonstrated that foundation models exhibit robust capabilities for next-token prediction during their pre-training phase Touvron et al. (2023b, a). Yet, this predictive strength is frequently underutilized in evaluating IT data quality. In SelectIT, we adopt a similar idea to evaluate IT data. Specifically, we calculate the next-token (from 1 to $K$) probability based on the rating prompt $RP$ and sample $S$. The score token with the highest probability is then considered as the sample’s quality.
$S^{base}=\underset{k\in\{1,\ldots,K\}}{\arg\max}\,P^{\prime}_{k},P^{\prime}_{k}=\left(\frac{e^{P_{k}}}{\sum_{j=1}^{K}e^{P_{j}}}\right)$ (2)
where $P_{k}$ and $P^{\prime}_{k}$ mean the probability and softmaxed probability of token $k$. $K$ means the number of scores to be considered.
The probability distribution among score tokens reflects the internal uncertainty of LLMs on sample evaluation. The higher $P^{\prime}_{S^{base}}$, the more confidence of LLMs, which is not well exploited in Equation 2. To capture this subtle difference, we introduce the token-level self-reflection (Token-R), which uses the distribution between tokens that reflect the internal uncertainty of LLMs, to enhance the credibility of quality assessment. Specifically, we assess the average disparity between the predicted $S^{base}$ token and the other, where the greater the disparity, the more the confidence of LLMs. This disparity is then utilized to refine the original $S^{base}$, resulting in a token-level score $S^{token}$.
$S^{token}=S^{base}\times\underbrace{\frac{1}{K-1}\sum\limits_{i=1}^{K}|P^{\prime}_{i}-P^{\prime}_{S^{base}}|}_{Uncertainty}$ (3)
3.2 Sentence-level Self-Reflection
Different prompts can significantly affect outputs of LLMs Kung et al. (2023); Peng et al. (2023b), introducing uncertainty into IT data evaluation at the sentence level. To make better use of this uncertainty to bolster the reliability of our method, we implement sentence-level self-reflection (Sentence-R). Building upon Token-R, we devise $K$ semantically similar rating prompts$\{RP_{0},RP_{1},\ldots,RP_{K}\}$ to obtain a series of quality scores $\{S^{token}_{0},S^{token}_{1},\ldots,S^{token}_{K}\}$ based on a given sample $S$. We calculate the average of these scores to represent the overall quality of sample $S$, because of the importance of incorporating assessments from diverse prompts. Additionally, we use the standard deviation to quantify the LLMs’ uncertainty to rating prompt; a higher standard deviation suggests greater sensitivity to prompt variation, while a lower standard deviation indicates more consistent and confident quality ratings by LLMs (Zhou et al., 2020). By integrating a holistic sample evaluation with the quantification of model uncertainty, we derive the sentence-level score $S^{sent}$, offering a more nuanced and reliable measure of IT data quality.
$S^{sent}=\frac{\textbf{Avg}\{S^{token}_{i}\}^{K}_{i=1}}{1+\alpha\times{\textbf{Std}\{S^{token}_{i}\}^{K}_{i=1}}}$ (4)
where ${\textbf{Avg}\{\cdot\}}$ and ${\textbf{Std}\{\cdot\}}$ respectively denote the mean and standard deviation of $S^{token}_{i}$, K means the number of rating prompts $RP$. Moreover, we use the uncertainty factor $\alpha$ to control for the impact of the uncertainty of LLMs on overall scores.
3.3 Model-level Self-Reflection
A sample affirmed by multiple foundation models can truly be deemed as high-quality. Different foundation models have different quality assessments of the sample, which introduce model-level uncertainty. To maximize the utilization of this uncertainty, we introduce model-level self-reflection (Model-R). This strategy is in harmony with our principles, as it leverages the capabilities of existing open-source models without the need for additional resources or the complexities associated with fine-tuning However, the challenge lies in the diverse capabilities of various LLMs and determining how to reasonably combine their sample evaluation based on their performance. It is widely acknowledged that the capabilities of LLMs tend to increase with their parameter count Hendrycks et al. (2021). Thus, we suggest using the parameter count of LLMs as an initial metric for assessing their capabilities to properly weight sample quality scores. Given $N$ foundation models with parameter counts $\{\theta_{1},\theta_{2},\ldots,\theta_{N}\}$ and their respective sentence-level scores for a sample $S$ being $\{S^{sent}_{0},S^{sent}_{1},\ldots,S^{sent}_{N}\}$, we formulate the model-level score $S^{model}$ to reflect a comprehensive evaluation of sample quality.
$Quality\propto{S^{model}}=\sum_{i=1}^{N}\left(\frac{\theta_{i}}{\sum_{j=1}^{N}\theta_{j}}\times S^{sent}_{i}\right)$ (5)
where $N$ means the number of the foundation models. We use $S^{model}$ as the final evaluation of sample $S$ in SelectIT. The higher $S^{model}$, the better sample quality. We sort the samples in descending order based on their $S^{model}$ and then select the top-ranked samples as high-quality data.
3.4 Selective Alpaca
We apply SelectIT to the widely-used Alpaca-GPT4 (Peng et al., 2023a). Specifically, we use the most popular LLaMA-2 (7B, 13B, 70B) as our foundation models and set the hyper-parameters $K=5$, which decides the range of LLMs rating in Token-R and the number of rating prompts in Sentence-R. We finally select the top 20% as the high-quality data and obtain a novel IT dataset called Selective Alpaca.
4 Experiments
ID   System External MMLU BBH GSM TydiQA CodeX AE Overall
Model Data AVG $\Delta$ (normal-$\uparrow$)
Base Model: LLaMA-2-7B Implemented Existing Method
 1    Alpaca-GPT4 46.2 39.2 15.0 43.3 27.8 33.7 34.3 -
 2    LIMA 45.8 38.5 14.0 45.5 24.2 33.1 33.5 -0.7
 3    1 + AlpaGasus 46.8 39.2 14.5 48.4 26.5 34.6 35.0 +0.8
 4    1 + Q2Q 46.7 39.8 16.5 45.5 28.1 35.1 35.3 +1.1
 5    1 + Instruction Mining 47.0 40.0 16.5 47.8 29.6 34.4 35.9 +1.7
[…]
 6    1 + Token-R 47.7 38.7 15.0 48.6 28.0 35.3 35.5 +1.3
 7    1 + Sentence-R 46.7 39.9 16.5 48.9 27.8 34.5 35.7 +1.5
 8    1 + Model-R 47.5 38.0 16.5 45.9 28.7 35.6 35.4 +1.2
[…]
 9    SelectIT (6 + 7 + 8) 47.7 40.8 17.5 47.8 29.8 35.4 36.5 +2.3
Base Model: LLaMA-2-13B Implemented Existing Method
10    Alpaca-GPT4 55.7 47.3 31.0 49.1 41.8 46.5 45.2 -
11    LIMA 54.6 48.3 29.5 51.2 34.4 43.6 43.6 -1.6
12   10 + AlpaGasus 54.1 49.3 32.0 52.6 39.3 47.5 45.8 +0.6
13   10 + Q2Q 55.3 48.5 34.0 50.8 42.3 47.7 46.4 +1.2
14   10 + Instruction Mining 55.5 49.7 33.0 51.2 41.5 46.1 46.2 +1.0
[…]
15   10 + Token-R 55.3 47.9 31.0 53.8 40.8 46.2 45.8 +0.6
16   10 + Sentence-R 55.2 48.5 34.0 54.2 41.5 47.3 46.8 +1.6
17   10 + Model-R 55.5 47.2 33.5 53.0 42.2 47.1 46.4 +1.2
[…]
18   SelectIT (15 + 16 + 17) 55.5 50.1 34.5 54.4 43.2 48.2 47.7 +2.5
Table 1: The overall results on IT. “Codex” and “AE” mean HumanEval and AlpacaEval benchmark respectively.
4.1 Setups
Benchmark
To gain a more comprehensive understanding of the capabilities of LLMs, we evaluate our approach in diverse downstream tasks Wang et al. (2023); Ivison et al. (2023). Factual knowledge: We use the Massive Multitask Language Understanding dataset (MMLU Hendrycks et al. (2021)) to assess the factual knowledge of LLMs and report 5-shot results. Reasoning: We evaluate the reasoning abilities of LLMs using two widely utilized datasets: the Grade School Math dataset (GSM (Cobbe et al., 2021)) and Big-Bench-Hard (BBH Suzgun et al. (2022)) with the CoT setting Wei et al. (2022b). Multilinguality: we assess this ability by TyDiQA, a multilingual question-answering benchmark that encompasses 11 diverse languages, with the gold-passage setup. Coding: We evaluate this ability using the HumanEval dataset (Chen et al., 2021) and report pass@10 results with a temperature of 0.8. Open-ended generation: We utilize AlpacaEval (Dubois et al., 2023), which employs GPT-4 to effectively assess model outputs. This can evaluate whether the text produced by LLMs aligns with humans.
Implementation Details
We use LLaMA-2 as our testbed. We fine-tune it for 3 epochs, with a batch size of 128. We use Adam with $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999, and the cosine learning rate scheduler starts from 2e--5, and decays to 0. we opted for a 4096 input length because it can show the best performance of LLMs. We employ the beam $=4$ for decoding. We set the temperature parameter to 0.8 and the top--p sampling parameter to 0.9 to improve the originality of the output text while ensuring the accuracy and relevance of the content.
Baselines
The compared baselines are:
  • Alpaca-GPT4 (Peng et al., 2023a) is a widely-used IT dataset that implements a self-instruct method to autonomously generate instructions by the advanced GPT4.
[…]
  • AlpaGasus (Chen et al., 2023a) involves utilizing the robust ChatGPT to score and select data from the original Alpaca dataset.
  • Q2Q (Li et al., 2023a) operates by training a precursor model, determining the quality of the IT data based on the two different loss values within this model.
  • Instruction Mining  (Cao et al., 2023) entails fitting data features and loss values to derive a formula for assessing data quality.
4.2 Main Results
We focus on the discussion of LLaMA-2-13B because both 7B and 13B models exhibit similar trends in Table 4. System (10) shows the vanilla IT on LLMs with the original Alpaca. By using the data selection strategies, the ability of LLMs has a moderate enhancement in Systems (12) to (14). Additionally, we can use $S^{base}$ as the input for Equations 4 and 5 to construct individual methods of Sentence-R and Model-R. Systems (15) to (17) illustrate that applying each submodule of SelectIT incrementally enhances LLMs’ performance, rivaling contemporary advanced methods.
Most remarkably, SelectIT markedly boosts LLaMA-2’s performance compared to vanilla IT in the System (18). This enhancement is particularly evident in the computational and reasoning tasks on the BBH and GSM benchmarks, compared to other IT data selection strategies. This may be attributed to the characteristics of selected data by SelectIT, and we will analyze this phenomenon in a later section. These gains in reasoning ability also positively impact the coding proficiency of LLMs. The improvement of LLMs on the TydiQA dataset is also obvious enough, which shows that SelectIT can effectively eliminate similar samples and retain sufficient diversity in multilingual aspects.
5 Analysis
We aim to answer the research questions through the following experiments:
  • How to select high-quality data in SelectIT?
  • Is SelectIT adaptable to various models and domains?
  • What are the advantages of Selective Alpaca?
5.1 Abalation Study of SelectIT
Effect of IT Data Quantity
Figure 3: Comparison of reasoning ability of LLMs with different proportions of Alpaca.
While SelectIT already excels at assessing and ranking samples effectively, selecting an appropriate number of samples in a redundant dataset remains a crucial aspect of our method. We divide the Alpaca dataset into multiple subsets ranging from 10% to 100% based on SelectIT’s evaluation and evaluate the reasoning ability of LLMs which is the SelectIT mainly enhances. As illustrated in Figure 3, compared to the full dataset, we observe that LLMs achieve optimal performance using only 20% of the data. Although the LLaMA-2-7B datasets of 20% and 40% displayed similar capabilities, the former required less fine-tuning time and computational resources. Hence, we opt for the 20% for implementing the SelectIT on the Alpaca dataset.
$K$ LLaMA-2-7B LLaMA-2-13B Overall
BBH GSM BBH GSM
3 40.1 17.5 49.6 34.0 35.3
5 40.8 17.5 50.1 34.5 35.7
7 41.2 18.0 49.5 34.0 35.7
9 40.5 17.0 50.3 34.0 35.5
Table 2: Effect of different $K$ on LLMs’ performance
ID Individual Method Unique (%) Overall (%)
6 Token-R 6.18 17.83
7 Sentence-R 40.81 63.98
8 Model-R 7.37 23.08
Table 3: The relationship between the SelectIT and the individual selection strategy. Sentence-R plays the most significant impact on the final rating of the IT data.
Effect of $K$
$K$ is a critical parameter for our method, impacting not only the range of scores assigned by the LLMs but also the number of rating prompts. We set $K={3,5,7,9}$ and apply the SelectIT algorithm for sample selection within the Alpaca to get different subset datasets. The results in Table 3 indicate that variations in the value of $K$ have a minimal impact on the overall performance of the LLMs. This is attributed to our multi-granularity self-reflection mechanism, which effectively enhances the accuracy and stability of sample selection. Although the model achieves competitive performance at $K$ values of 5 and 7, to minimize resource consumption, we set $K=5$ as the default value in SelectIT.
Datasets Datasets MMLU BBH GSM Tydiqa CodeX AE Overall
AVG $\Delta$ (normal-$\uparrow$)
LLaMA-2-7B Alpaca-GPT4 46.2 39.2 15.0 43.3 27.8 33.7 34.2 -
Selective Alpaca 47.7 40.8 17.5 47.8 29.8 35.4 36.5 +2.3
LLaMA-2-13B Alpaca-GPT4 55.7 47.3 31.0 49.1 41.8 46.5 45.2 -
Selective Alpaca 55.5 50.1 34.5 54.4 43.2 48.2 47.7 +2.5
Mistral-7B Alpaca-GPT4 52.5 51.7 33.5 51.1 54.7 43.1 47.8 -
Selective Alpaca 56.9 53.7 36.0 49.3 55.3 44.3 49.3 +1.5
Table 4: The overall results on IT of different LLMs.
[…]
COMET BLEU COMET BLEU COMET BLEU COMET BLEU COMET BLEU
SoTA Models
NLLB 54B 86.5 34.5 78.9 26.9 70.7 16.6 78.9 27.4 78.8 26.3
GPT-3.5 - 87.0 34.4 85.5 33.1 82.9 26.6 87.0 44.9 85.6 34.8
GPT-4 - 87.4 35.4 85.6 33.9 82.8 27.2 87.5 44.0 85.8 35.1
[…]
LLaMA-2 7B 76.4 19.0 82.7 30.4 75.0 18.2 71.8 17.0 76.5 21.1
TIM 7B 74.2 20.6 77.7 24.3 79.5 23.4 84.9 37.2 79.1 26.4
SWIE 7B 82.4 27.2 83.0 30.5 76.5 21.3 80.6 31.2 80.6 27.6
BigTranslate 13B 78.8 21.5 80.7 23.4 74.3 14.2 81.3 28.6 78.8 21.9
Bayling 13B 82.7 25.6 83.0 27.3 77.7 20.1 84.6 37.9 82.0 27.8
[…]
ALMA 7B 85.0 29.9 83.9 30.0 79.2 22.7 84.8 36.3 83.2 29.7
 w/ SelectIT 7B 85.2 30.2 84.1 30.4 80.0 24.2 85.3 37.3 83.7 30.5
[…]
 w/ SelectIT 13B 85.8 31.7 84.6 31.4 80.3 25.4 86.1 40.4 84.2 32.2
Table 5: The overall results on machine translation LLMs.
Effect of Different Reflection Strategy
We analyze the relationship between individual selection strategies and SelectIT, from the following two aspects. We first account for the number of high-quality data that only can be selected by a specific selection strategy, referred to as unique selection. Secondly, we calculate the percentage of the samples selected by individual selection strategies in Selective Alpaca, called overall selection. As shown in Table 3, Sentence-R plays the most important role in the final SelectIT strategy which also aligns with the model’s performance reported in Table 4, showing the rationality of our proposed uncertainty-aware self-reflection methods.
5.2 Robustness across Models and Domains
Various Foundation Models
Although Selective Alpaca achieved impressive improvements in LLaMA-2, applying it to other foundation models remains a challenging task. To address this, we apply Selective Alpaca on the Mistral-7B foundation model due to its superior performance Jiang et al. (2023) and present our results on the open-instruct benchmark alignment with the above test configuration. As depicted in Table 4, although Selective Alpaca is selected by the LLaMA-2 models, it is also applicable to the Mistral-7B and improves its capabilities across various tasks, especially on MMLU, BBH, and GSM benchmarks. This experiment fully demonstrates the flexibility of SelectIT which does not rely on a specific foundation model for data selection and the universality of Selective Alpaca which can effectively improve the capabilities of different series or scale LLMs.
Various Domain-specific Tasks
[…]
Table 5.1 shows that SelectIT consistently improves the translation performance of ALMA. Specifically, SelectIT mainly focuses on improving the translation from English to other languages, possibly because ALMA itself has strong English capabilities, which makes it difficult to improve. These results indicate that SelectIT is a versatile method that is effective for IT data selection and applicable to domain-specific tasks.
Method LLaMA-2 ALMA $\Delta$ (normal-$\uparrow$)
7B 13B 7B 13B
Full Dataset 34.3 45.2 29.7 31.5 -
 w/ Random (Full) 34.1 45.1 29.3 31.0 -1.2
 w/ Random (Unselected) 34.6 44.3 29.1 31.2 -1.5
 w/ Length 35.5 47.1 30.1 31.8 +3.8
 w/ SelectIT 36.5 47.7 30.5 32.2 +6.3
Table 6: Comparasion with selection variants.
Figure 4: Instruction embeddings representations of different selection strategies after T-SNE dimensionality reduction. The red and green points are representations of full Alpaca datasets and selected data respectively.
Figure 5: Left: The average length of samples. Right: The proportion of calculation type.
5.3 Insights of Selective Data Curation
Different Selection Strategies
This part compares three different selection strategies, namely, randomly selecting 20% in the full Alpaca, randomly selecting 20% in the unselected dataset, and selecting 20% data based on sample length Zhao et al. (2024), with our SelectIT in the IT and MT experiments. As shown in Table 6, the random-based strategies show certain performance degradation and the random selection in the unselected dataset is even worse, which reflects the effectiveness of our method from the side. Selection based on sample length is a simple approach to defining high-quality data, but it does not take into account the content of IT data, resulting in the limited performance of LLMs. SelectIT can significantly improve the abilities of LLMs.
Data Representation Analysis
This part investigates the relationship between the Selective Alpaca and the original datasets from a representation perspective. We follow Gao et al. (2024) to utilize the outputs of the last layer corresponding to the last token in the input sequence as sample representations. Subsequently, we employ T-SNE (Hinton and Roweis, 2002) for dimensionality reduction, mapping high-dimensional embeddings onto a 2D space. Figure 4 displays the intermediate representations generated by the full and Selective Alpaca respectively. It is observable that it is difficult for randomly selected data to distinguish abnormal data far from the center, making it challenging to define high-quality IT data. However, the majority of Selective Alpaca are concentrated around the central point, indicating that our dataset predominantly contains high-quality data near the center and effectively discards abnormal data distanced from the center which is a finding that strongly supports the conclusion of Table 6.
Data Characteristic Analysis
We analyze the Selective Alpaca from the following two perspectives, to explore why our dataset is better than the original dataset and its variants. Firstly, as shown in Figure  5, the length of instructions from the Selective Alpaca is significantly longer than those in the Alpaca dataset and AlpaGasus which is selected by ChatGPT. This implies that, with the same amount of data, our dataset contains more information, aligned with the results in Table 6. Secondly, by using ChatGPT to examine IT data types, we find a substantial increase in the proportion of computational problems in Selective Alpaca. This indicates that Selective Alpaca tends to select high-quality mathematical data, providing a solid explanation for the observed improvement in the reasoning abilities of LLMs as demonstrated in Table 4. Appendix A.1 shows the case study of comparing the Selective Alpaca with AlpaGasus.
6 Conclusion
This paper introduces a novel data selection strategy, SelectIT, for LLM instruction tuning, which uses LLM uncertainty to efficiently identify high-quality IT data without requiring additional resources. SelectIT includes three types of self-reflection: token, sentence, and model, which can individually and jointly improve the performance of IT data selection. By applying SelectIT to the Alpaca-GPT4 dataset, we introduce a compact and strong IT dataset, called Selective Alpaca. Different models and domain tasks demonstrate the effectiveness of SelectIT. Our analysis reveals that SelectIT effectively excludes abnormal data and tends to select longer and calculational data.
Limitation
This paper could be further strengthened as follows:
  • Instruction Data Quantity: Our findings suggest that prioritizing the top 20% of high-quality data optimizes results for Alpaca. Future studies might explore adjusting this threshold based on the data quality in different datasets to enhance performance. Additionally, it would be beneficial to further explore the incorporation of unselected data using curriculum learning (Liu et al., 2020; Zhan et al., 2021; Ding et al., 2021a).
  • Models at Different Scales: Our analysis is currently limited to models smaller than 30B parameters due to computational constraints. Investigating the efficacy of Selective Alpaca on larger-scale LLMs, ranging from 30B to 70B parameters, could provide valuable insights into the method’s scalability and applicability across various model sizes.
  • Diversity of Foundation Models: The study exclusively employs the LLaMA-2 as the foundational model for data selection. Broadening the scope to include alternative foundation models, such as Mistral, may reveal more nuanced data selection strategies and improve overall effectiveness.
  • Expansion to Additional Instruction Datasets: Although SelectIT has been applied to the Alpaca dataset due to its widespread adoption, extending this methodology to incorporate other instruction tuning datasets could offer substantial advantages to the broader LLM research community.
Appendix A Appendix
A.1 Case Study
As demonstrated in Table 6, we illustrate the selection tendencies of SelectIT in contrast to AlpaGasus, which leverages advanced ChatGPT for data selection. In samples 1 to 4, SelectIT shows a preference for instruction tuning data containing intricate mathematical problems that contribute to improving the reasoning skills of the LLMs. On the contrary, AlpacaGasus frequently chooses IT data in samples 5 to 7 that primarily offer solutions to queries or lack coherent reasoning, which might limit its effectiveness.
Figure 6: Examples of instruction tuning data selected by SelectIT or AlpaGasus.